# Beyond Text: Improving LLM’s Decision Making for Robot Navigation via Vocal Cues

### v1

@@ Abstract
This work highlights a critical shortcoming in text-based Large Language Models (LLMs) used for human-robot interaction, demonstrating that text alone as a conversation modality falls short in such applications. While LLMs excel in processing text in these human conversations, they struggle with the nuances of verbal instructions in scenarios like social navigation, where ambiguity and uncertainty can erode trust in robotic and other AI systems. We can address this shortcoming by moving beyond text and additionally focusing on the paralinguistic features of these audio responses. These features are the aspects of spoken communication that do not involve the literal wording (lexical content) but convey meaning and nuance through how something is said. We present Beyond Text; an approach that improves LLM decision-making by integrating audio transcription along with a subsection of these features, which focus on the affect and more relevant in human-robot conversations.This approach not only achieves a 70.26% winning rate, outperforming existing LLMs by 48.30%, but also enhances robustness against token manipulation adversarial attacks, highlighted by a 22.44% less decrease ratio than the text-only language model in winning rate. “Beyond Text” marks an advancement in social robot navigation and broader Human-Robot interactions, seamlessly integrating text-based guidance with human-audio-informed language models.
Machine Learning, ICML
@@ 1 Introduction
The rapid advancements in text-based Large Language Models (LLMs) like GPT-4 and the LLaMA series (Achiam et al., 2023; Touvron et al., 2023a, b) have significantly enhanced capabilities in collaborative human-robot interactions (HRI) especially in diverse settings, including task planning, or social navigation (Shah et al., 2023; Liu et al., 2023a; Li et al., 2023; Ren et al., 2023; Liang et al., 2023). Broadly speaking, for HRI tasks, we want to design the robot to think like a human and best carry out its tasks. We as humans interact and solve problems together, often relying on verbal communication as a key modality. If humans and robots interact to perform collaborative navigation task in large complex environments, human’s potential unfamiliarity with the space and spatial anxiety(Chan et al., 2012) leads to providing ambiguous or uncertain navigational guidance, and these uncertainty are not only in literal wording, but also vocal nuance (Prestopnik & Roskos-Ewoldsen, 2000; Golledge, 2003).While LLMs exhibit strong textual reasoning and perform adeptly in robotic tasks through textual prompt tuning, their ability to process and interpret audio information, particularly the nuanced vocal features in speech tones, remains limited. Although GPT-4 (Achiam et al., 2023) introduces an audio-to-text API, enhancing its capability to accept audio input, this feature primarily focuses on transcription rather than analyzing vocal features. This limitation becomes particularly salient in human audio-guided social navigation, where assessing the trustworthiness of human guidance is crucial for successful navigation (Dorbala et al., 2021; Francis et al., 2023; Firoozi et al., 2023).
@@
Moreover, to the best of our knowledge, there are no publicly available datasets that consider vocal cues for navigation tasks. To address this gap, we introduce Disfluent Navigational Instruction Audio Dataset (DNIA). This dataset contains audio clips that capture vocal and textual uncertainties, characteristic of real-life human speech in navigational contexts. DNIA focuses on disfluencies and nuanced vocal expressions, which are vital for enhancing LLMs’ ability to interpret the trustworthiness of human speech. We believe that these contributions are a significant first-step towards developing more natural and effective LLM-based robot-human interaction systems for social navigation.
@@
    Provided Evidence for Advancing LLM Navigation with Vocal Features. The current foundation model, GPT4, is unable to capture the nuanced human ambiguity via only textual input. We propose the integration of affective vocal cue analysis beyond text to let intelligent systems like robots synthesize these streams into actionable intelligence, and thus not only reason what the human said, but also how the human said. Our results show low bias and low variance: a 70.26%+ winning rate in detecting the uncertainty and generating appropriate next-step action and a low confidence score compared to only using LLMs to read textual transcription, indicating increased high confidence in analyzing human uncertainty. Beyond Text augments LLMs’ ability to model human uncertainty in navigational scenarios.
@@ 2 Related Works
Large Language Models for Robotics. Recent studies have demonstrated the versatility of LLMs in various robotic contexts. These include employing few-shot summarizing skills for task planning in robots (Wu et al., 2023; Liu et al., 2023a, c), interpreting contextual information for robotic locomotion decision-making (Shek et al., 2023), integrating visual feedback for open-world robotic manipulation (Lin et al., 2023a; Ahn et al., 2022), generating precise low-level commands for dynamic motion control systems (Wang et al., 2023; Chang et al., 2023a), and enhancing motion planning in autonomous driving (Mao et al., 2023). For human-robot interaction, Shah’s research (2023) suggests using LLMs for semantic prediction of object locations, thereby optimizing navigation. SayCan(2022) introduces multimodal input to generate the optimal long-horizon instructions and effectively score the affordance probability of the instruction for robotic task planning. Meanwhile, the KnowNo project (2023) develops a framework to assess and synchronize the uncertainty levels in LLM-based planning systems, enabling them to seek assistance when necessary. However, unlike our framework, KnowNo (2023) does not account for potentially unreliable human input but rather relies on human selection among generated options. Beyond Text uniquely concentrates on assessing human uncertainty through affective computing for robotic navigation, particularly when guided by human audio that might be ambiguous or incorrect. This approach allows robots to gauge the reliability of human-audio navigation instructions by analyzing vocal emotions, enabling more informed reasoning.
@@
Foundation Models with Audio. The ability to perceive and interpret auditory information is essential for AI agents in real world. GPT-4’s recent update(Achiam et al., 2023), incorporating Whisper’s audio-to-text feature (Radford et al., 2023), significantly advances foundational models’ capacity to transcribe audio inputs. However, this transcription process often overlooks nuanced vocal features inherent in speech. Parallel research efforts (Ghosal et al., 2023; Borsos et al., 2023) explore language models for generating audio from text and enhancing audio-visual understanding in videos. Notably, SALMONN (Tang et al., 2023) integrates speech audio encoders with a multimodal language model, enabling direct processing of audio inputs for tasks like speech translation, answering spoken queries etc. Despite these advancements, current research does not specifically address the use of audio inputs in aiding human-robot navigation or in extracting vocal uncertainty from speech.
@@
Uncertainty Modeling in LLMs. Gauging the trustworthiness of responses generated by LLMs remains an open challenge in today’s LLM era, with limited research on uncertainty quantification for natural language generation. Lin (2023b) compared multiple uncertainty measurements in the black-box language model, and Kuhn (2023) introduced entropy as a method to model uncertainty in the large language model.
In our approach, we take inspiration from prior works (Kuhn et al., 2023; Lin et al., 2023b) and define a confidence score $\mathcal{C}(\rho)$ that is inversely proportional to the Kullback-Leibler divergence between the current LLM’s prediction and ground truth human perception distribution to show the LLM’s confidence on uncertainty measurement.
3 Evaluation Metric: Confidence Measure
In this section, we first mathematically characterize the metric that could effectively capture the uncertainty in the decision-making of LLMs. As shown in Figure 2, we consider a setting where a robot asks the human for instructions, and then the human provides instructions in the form of audio instruction. This setting is adapted based on prior work on social navigation (Hu et al., 2019; Lakomkin et al., 2018; Eppe et al., 2016; Francis et al., 2023). In order to answer the question raised by the robot, the human replies with a voice prompt $v\in\mathcal{V}$ which belongs to set of audio responses $\mathcal{V}$, which is then converted to a textual description $\mathcal{S}$ using the best available transcription model $\mathcal{T}_{m}:\mathcal{V}\rightarrow\mathcal{S}$ (here $\mathcal{S}$ is the set of all possible sentences). The transcripted description $v$ acts as a prompt to the language model (denoted by $p_{\theta}$ with $\theta$ being the language model parameters), which outputs a response $y=[y_{1},y_{2},\cdots,y_{k}]$ by generating each token $y_{i}$ at a time $i$. The joint probability over the tokens can be written as $p_{\theta}(y|x)=\prod_{i=1}^{k}p_{\theta}(y_{i}|y_{1},y_{2}\cdots y_{k-1})$. To perform decision-making for robots, with a few-shot prompt that includes possible next steps in a few scenarios, the LLM generates candidate responses $\{y^{j}\}_{j=1}^{J}$ and chooses the most likely answer.
$\mathcal{C}(\rho)=\frac{1}{\text{KL}(\rho,\rho^{*})},$ (1)
where we note that the confidence measure will reduce as the KL divergence between $\rho$ and $\rho^{*}$ increases. Ideally, we would like to have the confidence measure as high as possible, which requires us to reduce $\text{KL}(\rho,\rho^{*})$. We first expand and note that
where $H(\rho)=-\sum_{j=1}^{J}\rho(y^{j})\log(\rho{(y^{j})})$ is the entropy of predicted distribution $\rho(y^{j})$ by the LLM which quantifies the amount of information contained in random variables. We note that since we don’t have any control over the second term in (2) because $\rho^{*}$ is unknown, we can focus on minimizing the first term in (2), which tells us to maximize the entropy in order to reduce the KL divergence between $\rho$ and $\rho^{*}$. We hypothesize that one of the major reasons for this uncertainty is the information loss while considering only textual information $x$ and losing various critical attributes of the input audio information $v$. From the standard information theory, one can quantify this loss in information as $\mathcal{I}_{l}=H(V)-I(X;V)$ as the difference between the entropy of the original audio input random variable $V$ and the mutual information between the text random variable $X$ and $V$ represented as $I(Y;V)$ which measures the amount of information that is shared between $X,V$. In this work, we empirically show that adding the vocal cue in the input helps to reduce the $\text{KL}(\rho,\rho^{*})$ by plotting the confidence measure (1) for robotic navigation tasks in Section 6.
@@ 4 Our Approach
In this section, we will explain our novel framework ‘Beyond Text’ that uses LLMs to model human trust based on audio-based navigational guidance. We have briefly described it in Figure 2 and will provide more details below.
4.1 Uncertainty Quantification
The process begins by converting the audio clips of navigational guidance into text using the open-source Whisper model (Radford et al., 2023), which we explain in detail in Appendix C. The transcribed text has two key functions for the LLM: identifying uncertainty-related semantic cues and formulating appropriate navigational steps. We focus on disfluency in human speech as an indicator of uncertainty, which affects the confidence measure of the LLMs (cf. 1). We categorize language uncertainty into three types as follows:
@@ 4.2 Vocal Affective Cue Analysis
To quantitatively measure pitch and loudness features, we design an algorithm (see Algorithm 1 in Appendix) to detect the max change of these two features in the audio clip. For the duration, we split the transcription sentence into sub-instructions and measured the duration of each sub-instructions. Force alignment, a process that synchronizes text fragments to their corresponding segments in an audio narration, is used to aligned the vocal features’ timestamps with the sub-instructions.
In this work, we focus on significant changes in these audio features that exceeded a predetermined threshold while excluding variations in the initial and final three seconds of recordings to avoid contamination by recording noise. This audio augmentation prompt serves as input for decision-making processes that leverage language models. In a way, we give the LLMs the ear to listen to human affective cues embedded inside the vocal clip, thus improving its reasoning skills.
@@ 4.3 In-Context Prompting
Beyond Text enables the language model to exploit its inherent reasoning capacities for strategic planning and decision-making tasks through In-Context Learning (ICL) (Yousefi et al., 2023; Liu et al., 2023a; Li et al., 2023), by providing only few-shot examples with no explicit training. We illustrate ICL (Yousefi et al., 2023) and Chain-of-Thought (Wei et al., 2022) by including three comprehensive examples that delineate all potential scenarios a robot might encounter while navigating human interactions for directional assistance. Specifically, we provide examples that: 1. only has language uncertainty; 2. only has vocal uncertainty; 3. has both semantic and vocal uncertainty signals. In the response phase, the model reason to suggest a spectrum of five potential actions. Notably, the first option (A) is always a paraphrased version of the transcription sentence without uncertainty, mimicking the scenario where the robot chooses to blindly follow the human audio-guided instruction; and the final option (E) is consistently presented as “ask another person nearby for direction,” ensuring a fail-safe in the decision-making process. Detailed prompts and LLM output examples are shown in the Appendix B.
@@ 5 New Dataset: Disfluent Navigational Instruction Audio Dataset (DNIA)
DNIA is a dataset comprising a diverse collection of human audio recordings. These recordings capture a spectrum of navigational disfluencies, including 500 audio clips contributed by twenty individuals, including ten males and ten females, and encompassing a wide array of native and non-native speaker vocal styles. Each clip, ranging from 10 to 30 seconds, includes a series of English directives necessary for navigating to a specified location. The audio files are in WAV format, normalized with a headroom of -2dB, facilitating ease of use and consistent audio quality.
We identified and categorized two types of disfluency within our dataset: language uncertainty (LU): 285 data clips and text correct by vocal tone uncertainty (VU): 215 data clips. The LU category is the audio that contains semantic disfluency signals, such as hesitation and language uncertainty. The VU category is particularly noteworthy, as it encompasses instances where the textual transcription alone does not manifest any indicators of uncertainty, but the vocal tone shows significant uncertainty clues.
To mimic real-world Human-Robot Interaction (HRI) scenarios as closely as possible, we align with the methodologies suggested in LLM evaluation literature (Bang et al., 2023; Chang et al., 2023b). We employ human annotation as one of the primary assessment techniques. This involves manual evaluation, which is particularly relevant in HRI contexts due to its ability to capture the nuanced complexities of human interaction. We use human evaluation to calculate the winning rate of analyzing human navigational guidance uncertainty in Section 6.3. Our approach entails labeling human audio-guided instructions through a selection process from a set of five multiple-choice options designed to represent the uncertainty inherent in the content accurately. We recruit a different set of 20 participants with diverse genders and dialect backgrounds to ensure a broad range of perspectives. We developed a user study website (see Figure 8) where these participants listened to navigational audio clips. The instructions emphasized careful listening and unbiased selection, with a constraint of a maximum of one-minute per clip to simulate real-world decision-making time pressures.

### v2

# Beyond Text: Utilizing Vocal Cues to Improve Decision Making in LLMs for Robot Navigation Tasks

@@ Abstract
While LLMs excel in processing text in these human conversations, they struggle with the nuances of verbal instructions in scenarios like social navigation, where ambiguity and uncertainty can erode trust in robotic and other AI systems. We can address this shortcoming by moving beyond text and additionally focusing on the paralinguistic features of these audio responses. These features are the aspects of spoken communication that do not involve the literal wording (lexical content) but convey meaning and nuance through how something is said. We present Beyond Text; an approach that improves LLM decision-making by integrating audio transcription along with a subsection of these features, which focus on the affect and more relevant in human-robot conversations.This approach not only achieves a 70.26% winning rate, outperforming existing LLMs by 22.16% to 48.30% (gemini-1.5-pro and gpt-3.5 respectively), but also enhances robustness against token manipulation adversarial attacks, highlighted by a 22.44% less decrease ratio than the text-only language model in winning rate. “Beyond Text” marks an advancement in social robot navigation and broader Human-Robot interactions, seamlessly integrating text-based guidance with human-audio-informed language models.
@@ 1 Introduction
The rapid advancements in text-based Large Language Models (LLMs) like GPT-4 and the Gemini series (Achiam et al., 2023; Reid et al., 2024) have significantly enhanced capabilities in collaborative human-robot interactions (HRI) especially in diverse settings, including task planning, or social navigation (Shah et al., 2023; Liu et al., 2023a; Li et al., 2023; Ren et al., 2023; Liang et al., 2023). Broadly speaking, for HRI tasks, we want to design the robot to think like a human and best carry out its tasks. We as humans interact and solve problems together, often relying on verbal communication as a key modality. If humans and robots interact to perform collaborative navigation tasks in large complex environments, human’s potential unfamiliarity with the space and spatial anxiety(Chan et al., 2012) leads to providing ambiguous or uncertain navigational guidance, and these uncertainties are not only in literal wording but also vocal nuance (Prestopnik & Roskos-Ewoldsen, 2000; Golledge, 2003). While LLMs exhibit strong textual reasoning and perform adeptly in robotic tasks through textual prompt tuning, their ability to process and interpret audio information, particularly the nuanced vocal features in speech tones, remains limited (example shown in Figure 1). Although GPT-4 Achiam et al. (2023) introduces an audio-to-text API, enhancing its capability to accept audio input, this feature primarily focuses on transcription rather than analyzing vocal features. This limitation becomes particularly salient in human audio-guided social navigation, where assessing the trustworthiness of human guidance is crucial for successful navigation (Dorbala et al., 2021; Francis et al., 2023; Firoozi et al., 2023).
@@
Moreover, to the best of our knowledge, existing available audio-guided navigation datasets, like RxR (Ku et al., 2020), do not consider vocal cues for navigation tasks. To address this gap, we introduce Disfluent Navigational Instruction Audio Dataset (DNIA). This dataset contains audio clips that capture vocal and textual uncertainties, characteristic of real-life human speech in navigational contexts. DNIA focuses on disfluencies and nuanced vocal expressions, which are vital for enhancing LLMs’ ability to interpret the trustworthiness of human speech. We believe that these contributions are a significant first step towards developing more natural and effective LLM-based robot-human interaction systems for social navigation.
@@
    Provided Evidence for Advancing LLM Navigation with Vocal Features. The current foundation models are unable to capture the nuanced human ambiguity via only textual input. We propose the integration of affective vocal cue analysis beyond text to let intelligent systems like robots synthesize these streams into actionable intelligence, and thus not only reason what the human said, but also how the human said. Our results show low bias and low variance: a 70.26%+ winning rate in detecting the uncertainty and generating appropriate next-step action and a low confidence score compared to only using LLMs to read textual transcription, indicating increased high confidence in analyzing human uncertainty. Notably, all three different large language models we tested, including GPT3.5, GPT3, and Gemini Pro, exhibited a significant increase in winning rates. Beyond Text augments LLMs’ ability to model human uncertainty in navigational scenarios.
@@ 2 Related Works
Large Language Models for Robotics. Recent studies have demonstrated the versatility of LLMs in various robotic contexts. These include employing few-shot summarizing skills for task planning in robots (Wu et al., 2023; Liu et al., 2023b), interpreting contextual information for robotic locomotion decision-making (Shek et al., 2023), integrating visual feedback for open-world robotic manipulation (Lin et al., 2023a; Ahn et al., 2022), and enhancing motion planning in autonomous driving (Mao et al., 2023). For human-robot interaction, Shah’s research Shah et al. (2023) suggests using LLMs for semantic prediction of object locations, thereby optimizing navigation. SayCan Ahn et al. (2022) introduces multimodal input to generate the optimal long-horizon instructions and effectively score the affordance probability of the instruction for robotic task planning. Meanwhile, Ren et al. (2023) develops a framework to assess and synchronize the uncertainty levels in LLM-based planning systems, enabling them to seek assistance when necessary. However, unlike our framework, KnowNo Ren et al. (2023) does not account for potentially unreliable human input but rather relies on human selection among generated options. Beyond Text uniquely concentrates on assessing human uncertainty through affective computing for robotic navigation, particularly when guided by human audio that might be ambiguous or incorrect. This approach allows robots to gauge the reliability of human-audio navigation instructions by analyzing vocal emotions, enabling more informed reasoning.
@@
Foundation Models with Audio. The ability to perceive and interpret auditory information is essential for AI agents in the real world. GPT-4’s recent updateAchiam et al. (2023), incorporating Whisper’s audio-to-text feature (Radford et al., 2023), significantly advances foundational models’ capacity to transcribe audio inputs. However, this transcription process often overlooks nuanced vocal features inherent in speech. Parallel research efforts (Ghosal et al., 2023; Borsos et al., 2023) explore language models for generating audio from text and enhancing audio-visual understanding in videos. Notably, SALMONN Tang et al. (2023) integrates speech audio encoders with a multimodal language model, enabling direct processing of audio inputs for tasks like speech translation, answering spoken queries, etc. Despite these advancements, current research does not specifically address the use of audio inputs in aiding human-robot navigation or in extracting vocal uncertainty from speech.
@@
Uncertainty Modeling in LLMs. Gauging the trustworthiness of responses generated by LLMs remains an open challenge in today’s LLM era, with limited research on uncertainty quantification for natural language generation. Lin Lin et al. (2023b) compared multiple uncertainty measurements in the black-box language model, and Kuhn Kuhn et al. (2023) introduced entropy as a method to model uncertainty in the large language model. In our approach, we take inspiration from prior works (Kuhn et al., 2023; Lin et al., 2023b) and define a confidence score $\mathcal{C}(\rho)$ that is inversely proportional to the Kullback-Leibler divergence between the current LLM’s prediction and ground truth human perception distribution to show the LLM’s confidence on uncertainty measurement.
Human Affective Analysis in Robotics. For social robots to effectively coexist and interact with humans, it is imperative that they comprehend human emotional states for decision-making processes. Emotion understanding from speech for human-robot interaction has been studied in (Lakomkin et al., 2018a), while deep reinforcement learning methods (Dorbala et al., 2021) and cognition model (Eppe et al., 2016) has been used to understand textual ambiguities in natural languages. RxR dataset (Ku et al., 2020) is a well-known audio dataset for visual language navigation but does not cover the affective cues embedded in the vocal instructions, which hold substantial information that can reveal human uncertainty and trust. Beyond Text introduces an audio-augmentation technique for large language models (LLMs) that analyzes textual and vocal cues in human audio commands. This approach enables more human-like robotic systems for social navigation by identifying emotions and ambiguities through both speech content and vocal tones.
@@ 3 Our Approach
In this section, we first mathematically characterize our novel framework, ‘Beyond Text,’ which uses LLMs to model human trust and uncertainty based on audio-based navigational guidance. As shown in Figure 2, we consider a setting where a robot asks the human for instructions, and then the human provides instructions in the form of audio instruction. This setting is adapted based on prior work on social navigation (Hu et al., 2019; Lakomkin et al., 2018b; Eppe et al., 2016; Francis et al., 2023). In order to answer the question raised by the robot, the human replies with a voice prompt $v\in\mathcal{V}$, which belongs to set of audio responses $\mathcal{V}$, which is then converted to a textual description $\mathcal{S}$ using the best available transcription model $\mathcal{T}_{m}:\mathcal{V}\rightarrow\mathcal{S}$ (here $\mathcal{S}$ is the set of all possible sentences). Simultaneously, $v$ is mapped to an affective cue set $\mathcal{K}$ via an affective cue model $\mathcal{AC}:\mathcal{V}\rightarrow\mathcal{K}$. The combination of textual and affective cues is denoted as:
$\mathcal{Q}(\mathcal{V})=\mathcal{S}\oplus\mathcal{K}$ (1)
which constitutes the prompt for the LLM (denoted by $p_{\theta}$ with $\theta$ being the language model parameters), which outputs a response $y=[y_{1},y_{2},\cdots,y_{k}]$ by generating each token $y_{i}$ at a time $i$. The joint probability over the tokens can be written as:
$p_{\theta}(y|\mathcal{Q}(v)=\prod_{i=1}^{k}p_{\theta}(y_{i}|y_{1},y_{2}\cdots y_{k-1})$ (2)
To perform decision-making for robots, with a few-shot prompt that includes possible next steps in a few scenarios, the LLM generates candidate responses $\{y^{j}\}_{j=1}^{J}$ and chooses the most likely answer. We have briefly described it in Figure 2 and will provide more details below.
3.1 Semantic Uncertainty Quantification
The process begins by converting the audio clips of navigational guidance into text using the open-source Whisper model (Radford et al., 2023), which we explain in detail in Appendix B. The transcribed text has two key functions for the LLM: identifying uncertainty-related semantic cues and formulating appropriate navigational steps. We focus on disfluency in human speech as an indicator of uncertainty, which affects the confidence measure of the LLMs (cf. 3). We categorize language uncertainty into three types as follows:
@@ 3.2 Vocal Affective Cue Analysis
To quantitatively measure pitch and loudness features, we design a lightweight yet effective algorithm (see Algorithm 1 in AppendixF) to detect the max change of these two features in the audio clip. For the duration, we split the transcription sentence into sub-instructions and measured the duration of each sub-instructions. We only mark the sub-instructions that are significantly longer than other sub-instructions in the same audio clips as uncertainty signals. Force alignment, a process that synchronizes text fragments to their corresponding segments in audio narration, is used to align the vocal features’ timestamps with the sub-instructions. Examples are shown in Figure 2 and detailed in Appendix C.
In this work, we focus on significant changes in these audio features that exceeded a predetermined threshold while excluding variations in the initial and final three seconds of recordings to avoid contamination by recording noise. This audio augmentation prompt serves as input for decision-making processes that leverage language models. In a way, we give the LLMs the ear to listen to human affective cues embedded inside the vocal clip, thus improving their reasoning skills.
@@ 3.3 In-Context Prompting
Beyond Text enables the language model to exploit its inherent reasoning capacities for strategic planning and decision-making tasks through In-Context Learning (ICL) (Yousefi et al., 2023; Liu et al., 2023a; Li et al., 2023), by providing only few-shot examples with no explicit training. We illustrate ICL (Yousefi et al., 2023) and Chain-of-Thought (Wei et al., 2022) by including three comprehensive examples that delineate all potential scenarios a robot might encounter while navigating human interactions for directional assistance. Specifically, we provide examples that: 1. only has language uncertainty; 2. only has vocal uncertainty; 3. has both semantic and vocal uncertainty signals. In the response phase, the model reason to suggest a spectrum of five potential actions. Notably, the first option (A) is always a paraphrased version of the transcription sentence without uncertainty, mimicking the scenario where the robot chooses to blindly follow the human audio-guided instruction; and the final option (E) is consistently presented as “ask another person nearby for direction,” ensuring a fail-safe in the decision-making process. Detailed prompts and LLM output examples are shown in the Appendix C.
@@ 3.4 Scoring Choices by Polling LLMs
Figure 3: a) Average Confidence Score by context and audio type for samples where the language model picks the same choices as human perception. The skinny bar is the error bar, representing the standard deviation of the confidence score. Error bars give a general idea of how precise a measurement is or, conversely, how far from the reported value the true (error-free) value might be. b) Average Confidence Score by context and audio type. An overall low variance is indicated by the increased average confidence.
@@ 4 New Dataset: Disfluent Navigational Instruction Audio Dataset (DNIA)
DNIA is a dataset comprising a diverse collection of human audio recordings. All the audio clips are grounded in the robot navigation settings, as we ask the volunteers to give instructions to a location inside the buildings they have been to. These recordings capture a spectrum of navigational disfluencies, including 500 audio clips contributed by twenty individuals, including ten males and ten females, and encompassing a wide array of native and non-native speaker vocal styles. Each clip, ranging from 10 to 30 seconds, includes a series of English directives necessary for navigating to a specified location. The audio files are in WAV format, normalized with a headroom of -2dB, facilitating ease of use and consistent audio quality.
We identified and categorized two types of disfluency within our dataset: language uncertainty (LU): 285 data clips and text correct by vocal tone uncertainty (VU): 215 data clips. The LU category is the audio that contains semantic disfluency signals, such as hesitation and language uncertainty. The VU category is particularly noteworthy, as it encompasses instances where the textual transcription alone does not manifest any indicators of uncertainty, but the vocal tone shows significant uncertainty clues. Regarding ethics and privacy concerns, our dataset is collected by a group of fellow students with written consent to anonymously publish the recordings of their voices for research purposes. We reviewed all the audio files to ensure no personal information is present, and the annotations (including audio filenames) can not be matched to a specific participant. Authors have human subjects training certificates from the university. Therefore, we follow the existing guidelines to ensure the DNIA dataset does not have any ethics issues, and we will publish the dataset soon after the review process for research purposes. Table 1 shows examples for the audio prompt in DNIA, more details can be found in Appendix E.
To mimic real-world Human-Robot Interaction (HRI) scenarios as closely as possible, we align with the methodologies suggested in LLM evaluation literature (Bang et al., 2023; Chang et al., 2023). We employ human annotation as one of the primary assessment techniques. This involves manual evaluation, which is particularly relevant in HRI contexts due to its ability to capture the nuanced complexities of human interaction. We use human evaluation to calculate the winning rate of analyzing human navigational guidance uncertainty in Section 5.4. Our approach entails labeling human audio-guided instructions through a selection process from a set of five multiple-choice options designed to represent the uncertainty inherent in the content accurately. We recruit a different set of 20 participants with diverse genders and dialect backgrounds to ensure a broad range of perspectives. We developed a user study website (see Appendix G) where these participants listened to navigational audio clips. The instructions emphasized careful listening and unbiased selection, with a constraint of a maximum of one minute per clip to simulate real-world decision-making time pressures.